a rectangle is one fourth as wide as it is long. it is 20 feel long. what is its perimiter?
Answer by Alan3354(69443)

a rectangle is one fourth as wide as it is long. it is 20 feel long. what is its perimiter?
------------------
Width = 1/4 x 20 = 5 feet
Perimeter = 2*Length + 2*Width
P = 2*20 + 2*5
P = 50 feet